Solve for the indicated variable if the line through the two given points has the given slope.
(a, 6) and (2, 8), m = −1 A= ?

Answer by stanbon(75887)

Solve for the indicated variable if the line through the two given points has the given slope.
(a, 6) and (2, 8), m = −1 A= ?
-----
slope = (8-6)/(2-a) = -1
---
2/(2-a) = -1
2 = a-2
a = 4
-----------
Cheers,
Stan H.